HW #3 for Chem 318 Due: April 29, 2019 1. Solute A has a K 2.5 for an extraction between water (phase 1) and chloroform (phas need help with number 2
Answer #1

Volume of stationary phase, Vs = 13.8 ml

Volume of mobile phase, Vm = 18.2 ml

Retention volume of solute, Vr = 85.5 ml

Partition coefficient/distribution coefficient, K = Cs/Cm = number of moles/ volume in stationary phase/number of moles/Volume in mobile phase = (retention volume of solute - volume of mobile phase)/volume of stationary phase =

= (Vr - Vm)/Vs = 85.5 ml - 18.2 ml/13.8 ml = 67.3/13.8 = 4.88

Capacity factor, k' =

K = k' (Vm/Vs) = 4.88

k' (18.2/13.8) = 4.88

k' = 67.344/18.2 = 3.7

Capacity factor, k' = 3.7